OpenStudy (anonymous):

prove: cosec θ + tan θ sec θ = cosec θ sec^2 θ Help...

OpenStudy (anonymous):

well sec is 1/cos and tan is sin/cos. cosec is 1/sin

OpenStudy (anonymous):

cosec=1/sin tan=sin/cos and sec=1/cos plug them in the equ

OpenStudy (anonymous):